OpenStudy (anonymous):

Can you help solve Integrate 2x/(x-5) using u-substitution

OpenStudy (kc_kennylau):

u=x-5

OpenStudy (dumbcow):

then x = u+5 and du =dx \[2 \int\limits \frac{u+5}{u} du = 2 \int\limits 1 + 5u^{-1} du\]

OpenStudy (anonymous):

I f you write it as follows it would be come easiers \[ \frac{2 x}{x-5}=\frac{10}{x-5}+2 \]

OpenStudy (anonymous):

\[2+\frac{ 5 }{ x-5 }+c\]

OpenStudy (anonymous):

so would this be the answer?

OpenStudy (kc_kennylau):

You forgot to times 2 as well as to integrate \(5u^{-1}\)